OpenStudy (anonymous):

Divide. Rewrite the expression so that the imaginary unit does not appear in the denominator. 5+8i/9i

OpenStudy (anonymous):

\[\frac{ 5+8i }{ 9i }\]

OpenStudy (anonymous):

multiply top and bottom by \(i\)

OpenStudy (anonymous):

\[\frac{ 5 + 8i }{ 9i }\] \[\frac{ 5 + 8i }{ 9i }\times \frac{ 9i }{ 9i }\] \[\frac{ 9i(5 + 8i) }{ -81 }\]

OpenStudy (anonymous):

get \[\frac{5i-8}{-9}\]

OpenStudy (anonymous):

then you can rewrite it as \[\frac{8}{9}-\frac{5}{9}i\]

OpenStudy (anonymous):

\[\frac{ 45i - 72 }{ -81 }\]

OpenStudy (anonymous):

\[\frac{ 8 }{ 9 } - \frac{ 5 }{ 9 }i\]